Allen bought 20 stamps at the post in 37 cents and 20 cents denominations. If the total cost of the stamps was $7.06, how many 37 cents did Allen buy?
Answer by elima(1433)

Allen bought 20 stamps at the post in 37 cents and 20 cents denominations. If the total cost of the stamps was $7.06, how many 37 cents did Allen buy?
==========================================================================
x+y=20
.37x+.20y=7.06
-------------------
y=20-x
.37x+.20(20-x)=7.06
.37x+4-.20x=7.06
.37x-.2x=7.06-4
.17x=3.06
x=18
18+y=20
y=20-18
y=2
Allen bought 18 37 cent stamps
:)